1. Consider the function f(x) = e-x defined on the interval 0 < x < 1. (a) Give an odd and an even extension of this function Please solve for part (b) and (c) thank you!
Answer #1

b) f(x) = e-*, 0<x<1 As the half range is 1, so L = 1 Fourier sine representation is given as F(x) 5 bn sin(nax) where n=1 bn

The Fourier cosine series is given as ao TX f(x) + Enean cos (bn 0) ao = 2 an COS nx n=1 Where ao = {l-f(x)dx = 2 se-* dx = -

e-* = F. (x) ) = 4ο + + Σ=1 α, cos nx 2 + (1 – (-1)e-1) coς ηπχ. (2) 1+12π2 c) Put x = O in (2), then 2 (1 – (-1)*2-1) n=1 e

= Σε κοπό n=1 2 (1 – e-2) - 2e-2 +e-1 1+ η2π2 Cancelling e-7, from both sides and taking 2e-2 to the left side, we get 2 2e-2

for this nice and lengthy answer kindly give a thumbs up.